A rifle with a barrel length of 64 cm fires a 12 g bullet with a horizontal speed of 420 m/s . The bullet strikes a block of wood and penetrates to a depth of 14 cm .

A) What resistive force (assumed to be constant) does the wood exert on the bullet?

B) How long does it take the bullet to come to rest after entering the wood?
Answer #1

Here ,

L = 0.64 m

m = 0.012 Kg

initial speed ,u = 420 m/s

depth , d = 14 cm = 0.14 m

a) in the wood , let the resistive force is F

Using work energy theorum

F * d = 0.5 * m * v^2

F * 0.14 = 0.5 * 0.012 * 420^2

F = 7560 N

the resistive force is 7560 N

b)

let the time taken is t
using second law of motion

F * t= m * v

7560 * t = 0.012 * 420

t = 6.67 *10^-4 s

the time taken is 6.67 *10^-4 s